OpenStudy (anonymous):

1. Solve the equation for y. 4x 5y = 9

OpenStudy (anonymous):

-4x-5y=9 srry :(

OpenStudy (anonymous):

9-4x=5y (9-4x)/5=y

OpenStudy (rulnick):

y=(9-4x)/5 or y = 9/5-(4/5)x or y = 1.8 - 0.8x, your choice :)

OpenStudy (anonymous):

it has to be in fraction form

OpenStudy (anonymous):

Yeah so 9-4x)/5=y